OpenStudy (anonymous):

Morgan is buying a house for $314,000. She is financing $275,000 and obtained 25-year, fixed-rate mortgage with 5.875% interest rate. How much are er monthly payments? A) $24,272.20 B)$21,296.15 C)$1,999.18 D)$1,754.06

OpenStudy (tkhunny):

Reasonableness - No Interest: 275000 / (25*12) = 275000 / 300 = 917 -- Okay, it will have to be greater than that, but not an order of magnitude greater. I think this rules out A and B. Generally, i = 0.05875 -- The annual interest rate. j = i/12 = 0.004895833... -- The monthly interest rate. v = 1/(1+j) = 0.995128019 -- The monthly discount factor. After that, it's a matter of basic principles. \(275000 = Pmt\cdot (v + v^{2} + ... + v^{300}) = Pmt\cdot\dfrac{v-v^{301}}{1-v}\) Thus, \(Pmt = 275000\dfrac{1-v}{v-v^{301}} = 1750.88\) Well, it's not quite one of the answers, but a little rounding difference can cause that with a mortgage of this size. Keeping 6 decimal places throughout, j = 0.004896 v = 0.995128 Pmt: 1750.88 That didn't help much. Just 4 decimal places? j = 0.0049 v = 0.9951 Pmt: 1756.56 Whoops! Too far. Then 5 decimal places? j = 0.00490 v = 0.99512 Pmt: 1752.51 What I seem to be saying is that rounding is important.